Given a sample of Radium-226 having half-life of 4 days. Find the probability, a nucleus disintegrates after 2 half lives. 


 (A) 1 (B) 1/2 


 (C) 1.5 (D) 3/4 


Grade:11

2 Answers

Ahush
6 years ago
At half life n/2 sample remains and n/2 decays . Similarly after 2 half lives n/4 remains and 3n/4 decays . So Probality = (3n/4)/n=3/4
Akash Kalita
5 years ago
It